A factory has a conveyor​ belt, Conveyor​ 1, which brings the supplies into the factory. The conveyor belt rises 8 ft for every 4 ft it runs. The factory has another conveyor​ belt, Conveyor​ 2, which sends the products out of the factory. This conveyor belt descends 21 ft for every 7 ft it runs. Find the slope of each conveyor belt. Which conveyor belt is​ steeper?


1
Expert's answer
2021-10-18T16:14:35-0400

The slope of Conveyor 1 is "\tan \alpha=\frac {8} {4}=2",

"\alpha = \arctan(2) = 63.42\degree" .

2.The slope of Conveyor 2 is "\tan\beta =\frac {-21} {7} =- 3",

"\beta = \arctan(-3) \approx -71.56\degree".

It means the second conveyor belt is steeper.